OpenStudy (anonymous):

What is the solution of the system? Use substitution. y=x+6y=2x

OpenStudy (campbell_st):

doesn't make a lot of sense... and equation with 2 equal signs...

OpenStudy (anonymous):

y=x+6y ---(1) y=2x ---(2) sub (2) into(1) 2x=x+6(2x) 2x=x-12x x=0 sub x=0 into (2) y=0

OpenStudy (anonymous):

Yeah I no!!

OpenStudy (anonymous):

Well y=x+6y=2x. I guess you can imply y=2x x+6y <-- substitute for y x+6(2x) = 13x If 13x = 2x, then x = 0. If x=0, then y=0 too.